# Models of Set Theory: Extensions and Dead-ends

### v5

@@ 3. TALLER CONSERVATIVE ELEMENTARY EXTENSIONS
(a) (Keisler-Kunen [Ke-1], essentially). It is a theorem of $\mathrm{ZFC}+\Diamond_{\omega_{1}}$ that $\mathcal{M}$ can be elementarily extended to a rather branchless model.
(b) (Shelah [Sh]). It is a theorem of $\mathrm{ZFC}$ that $\mathcal{M}$ can be elementarily extended to a rather branchless model.
(a) $(\mathbb{P},\leq_{\mathbb{P}})$ is directed if any given pair of elements of $\mathbb{P}$ has an upper bound.
(b) A subset $F$ of $\mathbb{P}$ is a filter over $(\mathbb{P},\leq_{\mathbb{P}})$ if the sub-poset $(F,\leq_{P})$ is directed.
(c) A filter over $(\mathbb{P},\leq_{\mathbb{P}})$ is maximal if it cannot be properly extended to a filter over $(\mathbb{P},\leq_{\mathbb{P}})$. [^19] Thus a maximal filter can be described as a maximally compatible subset of $(\mathbb{P},\leq_{\mathbb{P}}).$
(d) A subset $C$ of $\mathbb{P}$ is cofinal in $(\mathbb{P},\leq_{\mathbb{P}})$ if $\forall x\in\mathbb{P}\ \exists y\in C\ y\leq_{\mathbb{P}}x.$
5.11. Definition. Suppose $s$ is an infinite set.
(a) $[s]^{<\omega}$ is the directed poset of finite subsets of $s,$ ordered by containment. Note that $[s]^{<\omega}$ is a directed set with no maximum element.
(b) $\mathrm{Fin}(s,2)$ is the poset of finite functions from $s$ into $\{0,1\}$, ordered by containment (where a function is viewed as a set of ordered pairs).
5.12. Example. Given an infinite set $s$, and $a\subseteq s,$ let $\chi_{a}:s\rightarrow 2$ be the characteristic function of $a,$ i.e., $\chi_{a}(x)=1$ iff $x\in a.$ Let $F_{a}=$ finite approximations to $\chi_{a}$. $F_{a}$ is a maximal filter of $\mathrm{Fin}(s,2).$
(a) Every $\mathcal{M}$-definable directed poset $(\mathbb{D},\leq_{\mathbb{D}})$ with no maximum element has a cofinal chain of length $\omega_{1}.$
(b) Given any $\mathcal{M}$-definable poset $\mathbb{P}$, and any maximal filter $F$ of $\mathbb{P}$, if $F$ has a cofinal chain of length $\omega_{1}$, then $F$ is coded in $\mathcal{M}$.
5.14. Remark. If $\tau=(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$ is a ranked tree, then each branch of $\tau$ is a maximal filter over $(\mathbb{T},\ \leq_{\mathbb{T}}).$ This makes it clear that every Rubin model is rather branchless. Also, a rather branchless model of $\mathrm{ZF}$ is rather classless. To see this consider the ranked tree $\tau_{\mathrm{class}}$ defined within a model $\mathcal{M}$ of $\mathrm{ZF}$ as follows: The nodes of $\tau_{\mathrm{class}}$ are ordered pairs $(s,\alpha)$, where $s\subseteq\mathrm{V}_{\alpha}$, the rank of $(s,\alpha)$ is $\alpha$ and $(s,\alpha)<(t,\beta)$ if $\alpha\in\beta$ and $s=t\cap\mathrm{V}_{\alpha}$. It is easy to see that $\mathcal{M}$ is rather classless iff every branch of $\tau_{\mathrm{class}}^{\mathcal{M}}$ is $\mathcal{M}$-definable. Hence we have the following chain of implications:
Rubin $\Rightarrow$ rather branchless $\Rightarrow$ rather classless.
5.15. Theorem. (Rubin [^20] The attribution to Rubin is informed by fact that the credit for [RS, Lemma 2.3] is explicitly given to Rubin (at the end of the introduction of the paper). [RS, Corollary 2.4]). It is a theorem of $\mathrm{ZFC}+\Diamond_{\omega_{1}}$ that if $\mathcal{M}$ is a countable structure in a countable language, then $\mathcal{M}$ has an elementary extension of cardinality $\aleph_{1}$ that is a Rubin model. [^21] The proof of [RS, Lemma 2.3] has a small gap (it is assumed that $T_{0}\vdash\exists\overline{x}\psi(\overline{x})$ instead of assuming the consistency of $T_{0}+\exists\overline{x}\psi(\overline{x})$). The improved proof was presented in [En-1, Theorem 2.1.3] and is reproduced here in the Appendix.
5.16. Definition. Suppose $\mathcal{M}$ is a model of $\mathrm{ZF}$. $\mathcal{M}$ is weakly Rubin if (a) and (b) below hold:
(a) $\mathcal{M}$ is rather classless (this is the asymptotic case of $(ii)$ below, if one could use $a=\mathrm{V}$).
(b) For every element $a$ of $\mathcal{M}$ that is infinite in the sense of $\mathcal{M}$ the following two statements hold:
$(i)$ $\left([a]^{<\omega}\right)^{\mathcal{M}}$ has a cofinal chain of length $\omega_{1}.$
$(ii)$ If $F$ is a maximal filter of $\mathrm{Fin}^{\mathcal{M}}(a,2)$ and $F$ has a cofinal chain of length $\omega_{1}$, then $F$ is coded in $\mathcal{M}$.
As demonstrated in the Appendix, Schmerl’s strategy of $\Diamond_{\omega_{1}}$-elimination in [Schm-2] (which is based on an absoluteness argument first presented by Shelah [Sh]) can be employed to build weakly Rubin models within ZFC. Thus we have:
5.17. Theorem (Rubin-Shelah-Schmerl) It is a theorem of $\mathrm{ZFC}$ that every countable model of $\mathrm{ZF}$ has an elementary extension to a weakly Rubin model of cardinality $\aleph_{1}$.
5.18. Theorem. Every countable model $\mathcal{M}_{0}\models\mathrm{ZF}$ has an elementary extension $\mathcal{M}$ of power $\aleph_{1}$ that has no proper end extension to a model $\mathcal{N}\models\mathrm{ZF}$. Thus every consistent extension of $\mathrm{ZF}$ has a model of power $\aleph_{1}$ that has no proper end extension to a model of $\mathrm{ZF}$.
Proof [^22] A more direct proof can be given using the machinery of recursive set functions, but in the interest of conceptual clarity, we opted for the proof presented here.. We begin with a basic fact that will be called upon in the proof.
@@
Proof. The assumptions on $\mathcal{M}$ and $\mathcal{N}$ readily imply that $\omega^{\mathcal{M}}=\omega^{\mathcal{N}}$ and $\mathrm{Ord}^{\mathcal{M}}\subseteq_{\mathrm{end}}\mathrm{Ord}^{\mathcal{N}}$. Recall that within $\mathrm{ZF}$, given any set parameter $a$ one can define the inner model $\mathrm{L}(a)$ as
$\bigcup\limits_{\alpha\in\mathrm{Ord}}\mathrm{L}_{\alpha}(a)$,
where $\mathrm{L}_{0}(a)$ is the transitive closure of $\{a\},$ $\mathrm{L}_{\alpha+1}(a)$ is the set of all subsets of $\mathrm{L}_{\alpha}(a)$ that are parametrically definable in $(\mathrm{L}_{\alpha+1}(a),\in),$ and for limit $\alpha,$
$\mathrm{L}_{\alpha}(a)=\bigcup\limits_{\beta\in\mathrm{\alpha}}\mathrm{L}_{\beta}(a).$
It is well-known that for each axiom $\varphi$ of $\mathrm{ZF}$, the statement $\forall a\ \varphi^{\mathrm{L}(a)}$ is provable in $\mathrm{ZF}$, and in particular, $\mathrm{ZF}$ can prove the following statement:
$\forall a\ \exists\alpha\in\mathrm{Ord\ }\left(\exists x\mathrm{\ }(x=\left[a\right]^{<\omega})\right)^{\mathrm{L}_{\alpha}(a)}$ (indeed $\alpha=2$ works).)).) .
On the other hand, the formula expressing $x=\mathrm{L}_{\alpha}(a)$ is provably $\Delta_{1}$ (in parameters $\alpha$ and $a$), so for a fixed $a\in M$ there is for some $\alpha\in\mathrm{Ord}^{\mathcal{M}}$ such that $\left(\left[a\right]^{<\omega}\right)^{\mathrm{L}_{\alpha}^{\mathcal{M}}(a)}=\left(\left[a\right]^{<\omega}\right)^{\mathrm{L}_{\alpha}^{\mathcal{N}}(a)}.$ Thus $\left(\left[a\right]^{<\omega}\right)^{\mathcal{M}}=\left(\left[a\right]^{<\omega}\right)^{\mathcal{N}}$ since
$\left(\left[a\right]^{<\omega}\right)^{\mathcal{M}}=\left(\left[a\right]^{<\omega}\right)^{\mathrm{L}_{\alpha}^{\mathcal{M}}(a)}$ and $\left(\left[a\right]^{<\omega}\right)^{\mathcal{N}}=\left(\left[a\right]^{<\omega}\right)^{\mathrm{L}_{\alpha}^{\mathcal{N}}(a)}.$
[…]
Given a countable model $\mathcal{M}_{0}$ of $\mathrm{ZFC}$, by Theorem 5.17 there is a weakly Rubin model $\mathcal{M}$ that elementary extends $\mathcal{M}_{0}$. By Theorem 5.1, to prove Theorem 5.18 it is sufficient to verify that every end extension $\mathcal{N}$ of $\mathcal{M}$ that satisfies $\mathrm{ZF}$ is a conservative extension. Towards this goal, suppose $\mathcal{M}\subsetneq_{\mathrm{end}}\mathcal{N}\models\mathrm{ZF}$. In light of Remark 2.6(e), it suffices to show that $\mathcal{M}\subsetneq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$.
To show $\mathcal{M}\subsetneq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$, suppose for $a\in M$ and $s\in N$, $\mathcal{N}\models s\subseteq a$ for some $a\in N.$ We will show that $s\in M.$ By Fact ()($\nabla$)( ∇ ) we may assume that $a$ is infinite as viewed from $\mathcal{M}$. Note that this implies that $\mathcal{M}$ views $[a]^{<\omega}$ as a directed set with no maximum element. Fact ()($\nabla$)( ∇ ), assures us that:
()($\ast$)( ∗ ) $\ \forall m\in M$ $\left([m]^{<\omega}\right)^{\mathcal{M}}=\left([m]^{<\omega}\right)^{\mathcal{N}},$ and
Since $\mathrm{Fin}(m,2)\subseteq[m\times\{0,1\}]^{<\omega}$, ()($\ast$)( ∗ ) implies:
[…]
Let $\chi_{s}\in N$ be the characteristic function of $s$ in $\mathcal{N}$, i.e., as viewed from $\mathcal{N}$, $\chi_{s}:a\rightarrow\{0,1\}$ and $\forall x\in a(x\in s\leftrightarrow\chi_{s}(x)=1)$. Let $F_{s}$
@@
[…]
$q_{\alpha}:=\mathrm{Ext}_{\mathcal{N}}(\chi_{s}\upharpoonright p_{\alpha})=\mathrm{Ext}_{\mathcal{M}}(\chi_{s}\upharpoonright p_{\alpha}).$
Therefore, by the assumption that $\mathcal{M}$ is weakly Rubin, $F_{s}$ is $\mathcal{M}$-definable and thus coded in $\mathcal{M}$ by some $m\in M$. This makes it clear that $s=\cup m$, this shows that $s\in$ $N$, which concludes the verification that $\mathcal{M}\subsetneq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}.$ As explained earlier, this is sufficient to establish Theorem 5.18. $\square$
@@
5.19. TheoremNo model of $\mathrm{KPR+\Pi}_{\infty}^{1}$-$\mathrm{DC}$ has a conservative proper end extension to a model of KPR.
@@
5.20. TheoremEvery countable model $\mathcal{M}_{0}\models\mathrm{KPR+\Pi}_{\infty}^{1}$-$\mathrm{DC}$ has an elementary extension $\mathcal{M}$ of power $\aleph_{1}$ that has no proper end extension to a model $\mathcal{N}\models\mathrm{KPR.}$ Thus every consistent extension of $\mathrm{KPR+\Pi}_{\infty}^{1}$-$\mathrm{DC}$ has a model of power $\aleph_{1}$ that has no proper end extension to model of $\mathrm{KPR.}$
5.21. QuestionIs there an $\omega$-standard model of $\mathrm{ZF(C)}$ that has no proper end extension to a model of $\mathrm{ZF(C)}$?
5.22. Remark. Weakly Rubin models are never $\omega$-standard (even though $\omega$-standard rather classless models exist in abundance). However, a slight variation of the existence proof of Rubin models shows that a Rubin model $\mathcal{M}$ can be arranged to have an $\aleph_{1}$-like $\omega^{\mathcal{M}}.$
@@ 6. ADDENDUM AND CORRIGENDUM TO [En-3]
Theorem 2.2 of [En-3] incorrectly stated (in the notation of the same paper, where Gothic letters are used for denoting structures) that if $\mathfrak{A}\models\mathrm{ZFC}$, $\mathfrak{A}\prec_{\mathrm{cons}}\mathfrak{B}$ and $\mathfrak{B}$ fixes each $a\in\omega^{\mathfrak{A}},$ then $\mathfrak{A}$ is cofinal in $\mathfrak{B}$. As shown in Theorem 3.1 of this paper, it is possible to arrange taller conservative elementary extensions of models of $\mathrm{ZFC}$, thus Theorem 2.2 of [En-3] is false as stated. However, by ‘slightly’ strengthening the condition:
$\mathfrak{B}$ fixes each $a\in\omega^{\mathfrak{A}},$
to
$\mathfrak{B}$ fixes $\omega^{\mathfrak{A}},$
one obtains a true statement, as indicated in Corollary 4.2 of this paper. As pointed out in Remark 3.9, the former condition holds as soon whenever $\mathfrak{A}\prec_{\mathrm{cons}}\mathfrak{B}\models\mathrm{ZF}$.
@@
  • Part (b) of Corollary 2.3 of [En-3] is correct as stated in [En-3]: if $\mathfrak{A}$ is a Rubin model of $\mathrm{ZFC}$ and $\mathfrak{A}\prec\mathfrak{B}$ that fixes every $a\in\omega^{\mathfrak{A}}$, then $\mathfrak{B}$ is a conservative extension of $\mathfrak{A}$. This is really what the proof of part (a) of Corollary 2.3 of [En-3] shows. However, the following correction is needed: towards the end of the proof of part (a) of Corollary 2.3 of [En-3], after defining the set $X$, it must simply be stated as follows: But $X=Y\cap A,$ where $Y$ is the collection of members $a\in A$ satisfying $[\varphi(b,a)\wedge a\in R_{\alpha}]$. Since $\mathfrak{B}$ is an end extension of $\mathfrak{C}$, $Y\in C,$ so by claim ()($\ast$)( ∗ ), $X$ is definable in $\mathfrak{A}$.
  • Corollary 2.4 of [En-3] is correct as stated since it does not rely on the full force of Theorem 2.2 of [En-3] and only relies on the fact that no elementary conservative extension of a model of $\mathrm{ZFC}$ is an end extension. The latter follows from Corollary 4.2 of this paper.
@@
A.1. Theorem (Rubin-Shelah-Schmerl) It is a theorem of $\mathrm{ZFC}$ that every countable model of $\mathrm{ZF}$ has an elementary extension to a weakly Rubin model of cardinality $\aleph_{1}$.
The proof of Theorem A.1 has two distinct stages: in the first stage we prove within $\mathrm{ZFC}+\Diamond_{\omega_{1}}$ that every countably infinite structure in a countable language has an elementary extension to a Rubin model of cardinality $\aleph_{1}$ (Theorem 5.15), and then in the second stage we use the result of the first stage together with a forcing-and-absoluteness argument to establish the theorem. [^23] To my knowledge it is not known whether Theorem 5.15 (existence of Rubin elementary extensions) can be proved in $\mathrm{ZFC}$ alone. However, as shown here, $\mathrm{ZFC}$ can prove that every countable model of $\mathrm{ZF}$ has an elementary extension to a weakly Rubin model. See also Remark A.7.
@@
Fix a $\Diamond_{\omega_{1}}$ sequence $\left\langle S_{\alpha}:\alpha<\omega_{1}\right\rangle$.  Given a countable language $\mathcal{L}$, and countably infinite $\mathcal{L}$-structure model $\mathcal{M}_{0}$, assume without loss of generality that $M_{0}=\alpha_{0}\in\omega_{1}.$ We plan to inductively build two sequences $\left\langle\mathcal{M}_{\alpha}:\alpha<\omega_{1}\right\rangle$, and $\left\langle\mathcal{O}_{\alpha}:\alpha<\omega_{1}\right\rangle.$ The first is a sequence of approximations to our final model $\mathcal{M}.$ The second sequence, on the other hand, keeps track of the increasing list of ‘obligations’ we need to abide by throughout the construction of the first sequence. More specifically, each $\mathcal{O}_{\alpha}$ will be of the form $\{\{V_{n}^{\alpha},\ W_{n}^{\alpha}\}:n\in\mathbb{\omega}\}$, where $\{V_{n}^{\alpha},\ W_{n}^{\alpha}\}$ is pair of disjoint subsets $M_{\alpha}$ that are inseparable in $\mathcal{M}_{\alpha}$ and should be kept inseparable in each $\mathcal{M}_{\beta}$, for all $\beta\geq\alpha.$ Here we say that two disjoint subsets $V$ and $W$ of a model $\mathcal{N}$ are inseparable in $\mathcal{N}$ if there is no $\mathcal{N}$-definable $X\subseteq N$ such that $V\subseteq X,$ and $W\cap X=\varnothing.\mathfrak{\vskip 6.0pt plus 2.0pt minus 2.0pt}$
@@
At stage $\alpha+1,$ where $\alpha$ is a limit ordinal, we have access to a model $\mathcal{M}_{\alpha}$ (where $M_{\alpha}\in\omega_{1}$), and a collection $\mathcal{O}_{\alpha}$ of inseparable pairs of subsets of $M_{\alpha}$. We now look at $S_{\alpha}$, and consider two cases: either $S_{\alpha}$ is parametrically undefinable in $\mathfrak{A}_{\alpha},$ or not. In the latter case we “do nothing” and define $\mathcal{M}_{\alpha+1}:=\mathcal{M}_{\alpha}$ and $\mathcal{O}_{\alpha+1}:=\mathcal{O}_{\alpha}.$ But if the former is true, we augment our list of obligations via:
$\mathcal{O}_{\alpha+1}:=\mathcal{O}_{\alpha}\cup\{\left\{S_{\alpha},\ M_{\alpha}\backslash S_{\alpha}\right\}\}.\mathfrak{\vskip 6.0pt plus 2.0pt minus2.0pt}$
Notice that if $S_{\alpha}$ is parametrically undefinable in $\mathcal{M}_{\alpha},$ then $\{S_{\alpha},\ M_{\alpha}\backslash S_{\alpha}\}$ is inseparable in $\mathcal{M}_{\alpha}.$ Then we use Lemma A.2 below to build an elementary extension $\mathcal{M}_{\alpha+1}$ of $\mathcal{M}_{\alpha}$ that satisfies the following three conditions:
@@
(2) For every $\mathcal{M}_{\alpha}$-definable directed set $\mathbb{D}$ with no last element, there is some $d^{\ast}\in\mathbb{D}^{\mathcal{M}_{\alpha+1}}$ such that for each $d\in\mathbb{D}^{\mathcal{M}_{\alpha}},$ $\mathcal{N}\models\left(d<_{\mathbb{D}_{n}}d^{\ast}\right).$
@@
We now verify that $\mathcal{M}$ is a Rubin model. Suppose, on the contrary, that for some $\mathcal{M}$-definable partial order $\mathbb{P}$, there is a maximal filter $F\subseteq\mathbb{P}$ that is not $\mathcal{M}$-definable, and $F$ contains a cofinal $\omega_{1}$-chain $E$. By a standard Löwenheim-Skolem argument there is some limit $\alpha<\omega_{1}$ such that
$(5)$ $\ \ (\mathcal{M}_{\alpha},S_{\alpha})\prec(\mathcal{M},S),$ where $S_{\alpha}=S\cap\alpha.$
@@
Since $S_{\alpha}$ is a countable subset of some $\mathcal{M}$-definable partial order $\mathbb{P}$, and $E$ is chain of length $\omega_{1}$ that is cofinal in $\mathbb{P}$, there is some $e\in\mathbb{P}$ such that $p<_{\mathbb{P}}e$ for each $p\in S_{\alpha}.$ Fix some $\beta<\omega_{1}$ such that $e\in M_{\beta}.$ We claim that the formula $\varphi(x)=(x<_{P}e)$ separates $S_{\alpha}$ and $\mathbb{P}^{M_{\alpha}}\backslash S_{\alpha}$ in $\mathcal{M}_{\alpha}.$ This is easy to see since if $q\in\mathbb{P}^{M_{\alpha}}\backslash S_{\alpha}$, then by (5), $S_{\alpha}$ is a maximal filter on $\mathbb{P}^{M_{\alpha}}$, and therefore there is some $p\in S_{\alpha}$ such that $\mathcal{M}_{\alpha}\models$$p$ and $q$ have no common upper bound”, which by (5) implies the same to be true in $\mathcal{M}$, in turn implying that $\lnot(q<_{\mathbb{P}}e),$ as desired. We have arrived at a contradiction since on one hand, based on (2) and (6), the formula $\varphi(x)$ witnesses the separability of $S_{\alpha}$ and $M_{\alpha}\backslash S_{\alpha}$ within $\mathcal{M},$ and on the other hand $\{S_{\alpha},\ M_{\alpha}\backslash S_{\alpha}\}\in\mathcal{O}$ by (5), and therefore (4) dictates that $S_{\alpha}$ and $M_{\alpha}\backslash S_{\alpha}$ are inseparable in $\mathcal{M}.$ Thus the proof of Theorem A.1 will be complete once we establish Lemma A.2 below.

### v6

@@ 2. $(2)$
Since $\mathrm{Fin}(m,2)\subseteq[m\times\{0,1\}]^{<\omega}$, $(\ast)$ implies:
@@
Let $\chi_{s}\in N$ be the characteristic function of $s$ in $\mathcal{N}$, i.e., as viewed from $\mathcal{N}$, $\chi_{s}:a\rightarrow\{0,1\}$ and $\forall x\in a(x\in s\leftrightarrow\chi_{s}(x)=1)$. Let
@@
As noted in Example 5.12, $F_{s}$ is a maximal filter over $\mathrm{Ext}_{\mathcal{N}}(\mathrm{Fin}^{\mathcal{N}}(a,2))$, so by $(\ast\ast)$ and the assumption that $\mathcal{N}$ end extends $\mathcal{M}$, $F_{s}$ is a maximal filter over $\mathrm{Ext}_{\mathcal{M}}(\mathrm{Fin}^{\mathcal{M}}(a,2))$. Note that directed set $\mathrm{Ext}_{\mathcal{M}}\left([a]^{<\omega}\right)^{\mathcal{M}}$ has a cofinal chain $\left\langle p_{\alpha}:\alpha\in\omega_{1}\right\rangle$ thanks to the assumption that $\mathcal{M}$ is weakly Rubin. Together with $(\ast)$, this readily implies that $F_{s}$ has a cofinal chain $\left\langle q_{\alpha}:\alpha\in\omega_{1}\right\rangle,$ where:
$q_{\alpha}:=(\chi_{s}\upharpoonright p_{\alpha})^{\mathcal{N}}=(\chi_{s}\upharpoonright p_{\alpha})^{\mathcal{M}}.$
Therefore, by the assumption that $\mathcal{M}$ is weakly Rubin, $F_{s}$ is $\mathcal{M}$-definable and thus coded in $\mathcal{M}$ by some $m\in M$. This makes it clear that $\mathcal{M}\models\chi_{s}=\cup m$, and thus $s\in$ $M$ since $s$ is $\Delta_{0}$-definable from $\chi_{s}$. This concludes the verification that $\mathcal{M}\subsetneq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$, which as explained earlier, is sufficient to establish Theorem 5.18. $\square$
@@
5.19. TheoremNo model of $\mathsf{KPR+\Pi}_{\infty}^{1}$-$\mathsf{DC}$ has a conservative proper end extension to a model of $\mathsf{KPR}$.
@@
5.20. TheoremEvery countable model $\mathcal{M}_{0}\models\mathsf{KPR}+\Pi_{\infty}^{1}$-$\mathsf{DC}$ has an elementary extension $\mathcal{M}$ of power $\aleph_{1}$ that has no proper end extension to a model $\mathcal{N}\models\mathsf{KPR}\mathrm{.}$ Thus every consistent extension of $\mathsf{KPR}+\Pi_{\infty}^{1}$-$\mathsf{DC}$ has a model of power $\aleph_{1}$ that has no proper end extension to model of $\mathsf{KPR}\mathrm{.}$
5.21. QuestionIs there an $\omega$-standard model of $\mathsf{ZF(C)}$ that has no proper end extension to a model of $\mathsf{ZF(C)}$?
5.22. Remark. Weakly Rubin models are never $\omega$-standard (even though $\omega$-standard rather classless models exist in abundance). However, a slight variation of the existence proof of Rubin models shows that a Rubin model $\mathcal{M}$ can be arranged to have an $\aleph_{1}$-like $\omega^{\mathcal{M}}.$ Note, however, that the answer to Question 5.21 is known to be in the negative if “end extension” is strengthened to “rank extension”; this follows from of [En-2, Theorem 1.5(b)], (which asserts that no $\aleph_{1}$-like model of $\mathsf{ZFC}$ that elementarily end extends a model all of whose ordinals are pointwise definable has a proper rank extension to a model of $\mathsf{ZFC}$. The proof of [En-2, Theorem 1.5(b)] together with part (b) of Theorem 4.4 makes it clear that the same goes for $\mathsf{ZF}$, i.e., there are $\omega$-standard models of $\mathsf{ZF}$ that have no proper rank extension to a model of $\mathsf{ZF}$.
@@ 6. ADDENDUM AND CORRIGENDUM TO [En-3]
Theorem 2.2 of [En-3] incorrectly stated (in the notation of the same paper, where Gothic letters are used for denoting structures) that if $\mathfrak{A}\models\mathsf{ZF}$, $\mathfrak{A}\prec_{\mathrm{cons}}\mathfrak{B}$ and $\mathfrak{B}$ fixes each $a\in\omega^{\mathfrak{A}},$ then $\mathfrak{A}$ is cofinal in $\mathfrak{B}$. As shown in Theorem 3.1 of this paper, it is possible to arrange taller conservative elementary extensions of models of $\mathsf{ZFC}$, thus Theorem 2.2 of [En-3] is false as stated. However, by assuming that ‘slightly’ strengthening the conditions:
$\mathfrak{A}\models\mathsf{ZF}$, and $\mathfrak{B}$ fixes each $a\in\omega^{\mathfrak{A}},$
to:
$\mathfrak{A}\models\mathsf{ZF}$, and $\mathfrak{B}$ fixes $\omega^{\mathfrak{A}},$
one obtains a true statement, as indicated in Corollary 4.2 of this paper. [^35] To my knowledge, there is only one place in the published literature in which Theorem 2.2 of [En-3] has been used, namely in a recent paper of Goldberg and Steel [GS]. In the discussion after Lemma 3.5 of the aforementioned paper, it is noted that Theorem 2.2 of [En-3] implies that an elementary embedding $\pi:\mathcal{M}\rightarrow\mathcal{N}$ (where $\mathcal{M}$ and $\mathcal{N}$ are inner models of $\mathsf{ZFC}$) has the property that $\pi^{-1}[S]\in M$ for each $S\in N$ (paraphrased in [GS] as $\pi$ is a close embedding of $\mathcal{M}$ into $\mathcal{N}$’) iff for each parametrically definable subset $D$ of $\mathcal{N}$, $\pi^{-1}[D]$ is parametrically definable in $\mathcal{M}$. In the context of [GS], the relevant models $\mathcal{M}$ and $\mathcal{N}$ are both $\omega$-models (they are inner models, thus well-founded), so Corollary 4.2 of this paper applies. As pointed out in Remark 3.9, $\mathfrak{B}$ fixes each $a\in\omega^{\mathfrak{A}}$ whenever $\mathfrak{A}\prec_{\mathrm{cons}}\mathfrak{B}\models\mathsf{ZF}$.
@@
  • Part (b) of Corollary 2.3 of [En-3] is correct as stated in [En-3]: if $\mathfrak{A}$ is a Rubin model of $\mathsf{ZFC}$ and $\mathfrak{A}\prec\mathfrak{B}$ that fixes every $a\in\omega^{\mathfrak{A}}$, then $\mathfrak{B}$ is a conservative extension of $\mathfrak{A}$. This is really what the proof of part (a) of Corollary 2.3 of [En-3] shows. However, the following correction is needed: towards the end of the proof of part (a) of Corollary 2.3 of [En-3], after defining the set $X$, it must simply be stated as follows: But $X=Y\cap A,$ where $Y$ is the collection of members $a\in A$ satisfying $\varphi(b,a)\wedge a\in R_{\alpha}]$. Since $\mathfrak{B}$ is an end extension of $\mathfrak{C}$, $Y\in C,$ so by claim $(\ast)$, $X$ is definable in $\mathfrak{A}$.
  • Corollary 2.4 of [En-3] is correct as stated since it does not rely on the full force of Theorem 2.2 of [En-3] and only relies on the fact that no elementary conservative extension of a model of $\mathsf{ZF}$ is an end extension. The latter follows from Corollary 4.2 of this paper.
@@
A.1. Theorem (Rubin-Shelah-Schmerl) It is a theorem of $\mathsf{ZFC}$ that every countable model of $\mathsf{ZF}$ has an elementary extension to a weakly Rubin model of cardinality $\aleph_{1}$.
The proof of Theorem A.1 has two distinct stages: in the first stage we prove within $\mathsf{ZFC}+\Diamond_{\omega_{1}}$ that every countably infinite structure in a countable language has an elementary extension to a Rubin model of cardinality $\aleph_{1}$ (Theorem 5.15), and then in the second stage we use the result of the first stage together with a forcing-and-absoluteness argument to establish the theorem. [^36] It appears to be unknown whether Theorem 5.15 (existence of Rubin elementary extensions) can be proved in $\mathsf{ZFC}$ alone. However, as shown here, $\mathsf{ZFC}$ can prove that every countable model of $\mathsf{ZF}$ has an elementary extension to a weakly Rubin model. See also Remark A.7.
@@
Fix a $\Diamond_{\omega_{1}}$ sequence $\left\langle S_{\alpha}:\alpha<\omega_{1}\right\rangle$. Given a countable language $\mathcal{L}$, and countably infinite $\mathcal{L}$-structure model $\mathcal{M}_{0}$, assume without loss of generality that $M_{0}=\omega.$ We plan to inductively build two sequences $\left\langle\mathcal{M}_{\alpha}:\alpha<\omega_{1}\right\rangle$, and $\left\langle\mathcal{O}_{\alpha}:\alpha<\omega_{1}\right\rangle.$ The first is a sequence of approximations to our final model $\mathcal{M}.$ The second sequence, on the other hand, keeps track of the increasing list of ‘obligations’ we need to abide by throughout the construction of the first sequence. More specifically, each $\mathcal{O}_{\alpha}$ will be of the form $\{\{V_{n}^{\alpha},\ W_{n}^{\alpha}\}:n\in\mathbb{\omega}\}$, where $\{V_{n}^{\alpha},\ W_{n}^{\alpha}\}$ is pair of disjoint subsets of $M_{\alpha}$ that are inseparable in $\mathcal{M}_{\alpha}$ and should be kept inseparable in each $\mathcal{M}_{\beta}$, for all $\beta>\alpha.$ Here we say that two disjoint subsets $V$ and $W$ of a model $\mathcal{N}$ are inseparable in $\mathcal{N}$ if there is no $\mathcal{N}$-definable $X\subseteq N$ such that $V\subseteq X,$ and $W\cap X=\varnothing.\mathfrak{\vskip 6.0pt plus 2.0pt minus 2.0pt}$
@@
At stage $\alpha+1,$ where $\alpha$ is a limit ordinal, we have access to a model $\mathcal{M}_{\alpha}$ (where $M_{\alpha}\in\omega_{1}$), and a collection $\mathcal{O}_{\alpha}$ of inseparable pairs of subsets of $M_{\alpha}$. We now look at $S_{\alpha}$, and consider two cases: either $S_{\alpha}$ is parametrically undefinable in $\mathcal{M}_{\alpha},$ or not. In the latter case we ‘do nothing’, and define $\mathcal{M}_{\alpha+1}:=\mathcal{M}_{\alpha}$ and $\mathcal{O}_{\alpha+1}:=\mathcal{O}_{\alpha}.$ But if the former is true, we augment our list of obligations via:
$\mathcal{O}_{\alpha+1}:=\mathcal{O}_{\alpha}\cup\{\left\{S_{\alpha},\ M_{\alpha}\backslash S_{\alpha}\right\}\}.$
Notice that if $S_{\alpha}$ is parametrically undefinable in $\mathcal{M}_{\alpha},$ then $S_{\alpha},\ M_{\alpha}\backslash S_{\alpha}\}$ is inseparable in $\mathcal{M}_{\alpha}.$ Then we use Lemma A.2 below to build an elementary extension $\mathcal{M}_{\alpha+1}$ of $\mathcal{M}_{\alpha}$ that satisfies the following three conditions:
@@
(2) For every $\mathcal{M}_{\alpha}$-definable directed set $\mathbb{D}$ with no last element, there is some $d^{\ast}\in\mathbb{D}^{\mathcal{M}_{\alpha+1}}$ such that for each $d\in\mathbb{D}^{\mathcal{M}_{\alpha}},$ $\mathcal{M}_{\alpha+1}\models\left(d<_{\mathbb{D}}d^{\ast}\right).$
@@
We now verify that $\mathcal{M}$ is a Rubin model. Suppose, on the contrary, that for some $\mathcal{M}$-definable partial order $\mathbb{P}$, there is a maximal filter $F\subseteq\mathbb{P}$ that is not $\mathcal{M}$-definable, and $F$ contains a cofinal $\omega_{1}$-chain $E$. By a standard Löwenheim-Skolem argument there is some limit $\alpha<\omega_{1}$ such that:
[…]
@@
[…]